OpenStudy (anonymous):

find the distance between the two points (6,7) and (6,1) @Jonathan1997

OpenStudy (anonymous):

NO! I am going to help her! Lol!! So are you ready?

OpenStudy (anonymous):

Distance formula: sqrt ( (1-7)^2 + (6-6)^2)) = sqrt (36 + 0) = sqrt (36) = 6

OpenStudy (anonymous):

HAHAHA JONATHAN! Yeah i'm ready

OpenStudy (anonymous):

FML!!!! *applause for Easy!*

OpenStudy (anonymous):

wait @Easyaspi314 how did you do that?

OpenStudy (anonymous):

NO!!! I know how to!! I will teach you!

OpenStudy (anonymous):

you take the square root (sqrt) of (the difference of the two y-values and square that + the difference of the two x-values and square that).

OpenStudy (anonymous):

Screw it!!! by!

OpenStudy (anonymous):

the two y-values are y = 1 and y = 7, so 1-7 = -6 and square -6 you get + 36. Same for the two x-values, both are 0 so 0 squared is 0

OpenStudy (anonymous):

So the square root (sqrt) of 36 = 6

OpenStudy (shamil98):

.